### v1

# Leveraging the Bias-Variance Tradeoff in Quantum Chemistry for Accurate Negative Singlet-Triplet Gap Predictions: A Case for Double-Hybrid DFT

Abstract
Molecules that violate Hund’s rule—having first excited singlet state (S1) below the triplet state (T1)—are rare yet promising as efficient light emitters. Their high-throughput identification demands exceptionally accurate excited-state modeling to minimize false positives and negatives. Benchmarking twelve S1–T1 energy gaps, we find that local variants of ADC(2) and CC2 deliver excellent accuracy and speed for screening medium-sized molecules. Notably, while double-hybrid DFT approximations (e.g., B2GP-PLYP and PBE-QIDH) exhibit high mean errors ($>100$ meV) despite very low standard deviations ($\approx 10$ meV), exploring their parameter space reveals that a configuration with 75% exchange and 55% correlation reduces the mean error to below $5$ meV—albeit with increased variance. Using this low-bias parameterization as an internal reference, we correct the systematic error while maintaining low variance, effectively combining the strengths of both low-bias and low-variance DFT parameterizations to enhance overall accuracy. Our findings suggest that low-variance DFT methods—often overlooked due to high bias—can serve as reliable tools for predictive modeling in first-principles molecular design.
1 Introduction
The energy criterion for organic light-emitting diodes (OLEDs) to operate via thermally activated delayed fluorescence (TADF) is that the energy of the first excited singlet state (S1) must be no more than $0.1$ eV higher than that of the triplet state (T1)[1, 2]. This energy criterion facilitates thermally aided reverse intersystem crossing (RISC), enabling the theoretical complete transfer of the T1 population to the emissive state, S1[3]. A derivative of the triangular-shaped molecule, heptazine (aka heptaazaphenalene, 7AP), with a central nitrogen (N) atom in an anti-aromatic 12-annulene framework, showed delayed fluorescence without thermal assistance, indicating a negative S1–T1 gap (STG)[4]. This mechanism paves the way for designing OLEDs that leverage exothermic delayed fluorescence from inverted singlet and triplet excited states (DFIST). Direct spectroscopic evidence of a negative STG ($-0.047\pm 0.007$ eV) was obtained for pentaazaphenalene (5AP) using anion photoelectron spectroscopy and fluorescence measurements [5]. Negative STG was also was also observed in the transient photoluminescence data of dialkylamine-substituted 5AP[6]. Computational studies have further suggested that azaphenalenes (APs) with other substitution patterns, as well as the boron (B) analog of 1AP, boraphenalene (1BP), also exhibit negative STGs[7, 8, 9, 10, 11, 12, 13, 14]. Non-alternant hydrocarbons and their substituted analogs have been shown to have the potential to exhibit negative STGs[15, 16, 17, 18, 19]. Notably, substituted analogs of the non-fused bicyclic hydrocarbon have demonstrated negative STGs, attributed to through-bond charge-transfer states[20].
A comprehensive search across $12,880$ structurally diverse small molecules revealed no exceptions to Hund’s rule, indicating that achieving the electronic structure criteria for ${\rm STG}<0$ requires nontrivial molecular structures[21]. Moreover, the typical magnitudes of negative STG are much smaller or similar to the average errors associated with popular quantum chemistry and density functional theory approximations (DFAs). Consequently, due to the rarity of molecules with negative STGs and the limitations of current quantum chemical approximations, accurately identifying molecules exhibiting ${\rm STG}<0$ via high-throughput screening—with minimal false-positive and false-negative predictions—remains challenging. The linear-response time-dependent density functional theory (LR-TD-DFT) formalism provides qualitative accuracy—i.e., gets the sign of the STG for DFIST candidates correct—only within the double-hybrid (dh) DFT framework containing many-body correlation via the second-order perturbation theory (MP2). With the right combination of ingredients, dh-DFT can reliably predict a wide range of molecular properties[22, 23, 24, 23]. Nevertheless, selecting suitable exchange-correlation (XC) functionals within the dh-DFT framework remains challenging, especially for novel problems lacking sufficient reference data.
We present a scheme that leverages the bias-variance tradeoff within a single DFT formalism. To achieve this, we evaluate various computational methods for predicting twelve STGs in triangular molecules, benchmarking them against previously reported theoretical best estimates[14]. We examine dh-DFAs exhibiting very low variance for STG predictions, and reparametrize them to minimize the bias. These optimal low-bias models are used as internal references to correct the predictions of the low-variance models. The internally referenced scaling approach is demonstrated by predicting STGs with low bias and low variance, resulting in excellent overall accuracy. We discuss the merits and drawbacks of this approach for its further applications.
2 Dataset and Computational Methods
We consider twelve systems (see Figure 1) using equilibrium geometries from Ref. [14], determined at the CCSD(T) level with the frozen core approximation and the cc-pVTZ basis set. Using these geometries, we performed single-point excited state calculations with various methods. In addition, we collected TBEs of the S1 and T1 energies from Ref. [14] to serve as reference values for benchmarking. As stated in Ref. [14], structure 11 in Figure 1 (with $D_{3\rm h}$ symmetry) corresponds to a saddle point on the potential energy surface at the MP2/6-311G($d$,$p$) level, whereas the true minimum geometry exhibits $C_{3\rm h}$ symmetry (structure 12). Similarly, a separate study[25] found that the high-symmetry forms of structures 6, 7, and 8 are not true minima and structure 2 is a very shallow potential well at the CCSD(T)/cc-pVTZ level[25]. Overall, four of the twelve structures are transition states; however, we include them in our benchmarking of STGs due to the availability of TBE-level results and the lack of alternative high-quality data.
Figure 1: Nitrogen- and boron-centered [12]-annulene systems reported in Ref. [14].
Laplace-transformed, density-fitted, local versions of correlated methods—CC2 [26] and ADC(2) [27]—offer attractive speedups for modeling excited states of medium-sized molecules. We denote these methods as L-CC2 and L-ADC(2), and performed calculations with the cc-pVDZ, aug-cc-pVDZ, cc-pVTZ, and aug-cc-pVTZ basis sets, using the corresponding JKFIT and MP2FIT auxiliary basis sets. L-CC2 and L-ADC(2) are the same methods denoted as LT-DF-CC2 and LT-DF-ADC(2) elsewhere. L-CC2 and L-ADC(2) calculations were conducted using Molpro (version 2015.1) [28]. L-CC2 has been show to deviate from the canonical variant CC2 by $\leq 0.05$ eV, irrespective of the character of the excited states, including that of charge-transfer states for which CC2 can have huge error [29]. The systems studied in this work do not have charger transfer characteristics, hence CC2 delivers good accuracy for STGs in agreement with TBEs[14]. This implementation has demonstrated excellent accuracy in modeling the excited states of BODIPY derivatives [30, 31]. For comparison, we collected S1 and T1 energies obtained using the linear-response coupled-cluster methods, EOM-CCSD and CC3, with the aug-cc-pVDZ basis set from Ref. [14]. Additionally, we performed ADC(2) calculations using the cc-pVDZ, cc-pVTZ, aug-cc-pVDZ, and aug-cc-pVTZ basis sets with the resolution-of-the-identity (RI) approximation [32, 33] as implemented in QChem (version 6.0.2) [34].
We conducted LR-TD-DFT calculations using the Tamm–Dancoff approximation (TDA). Although TDA may not predict oscillator strengths as accurately as the random-phase approximation (RPA), RPA predictions of triplet energies are sensitive to triplet stability [35]. Therefore, TDA is preferred for automated high-throughput calculations, and previous work has shown that its STG predictions for azaphenalines are more accurate than that of RPA[21]. The performance of various DFAs for predicting STGs in azaphenalines has been discussed in prior studies [36, 37, 38].
This study explores several DFAs spanning all levels of the Jacob’s ladder [39], namely, local density approximation (LDA): VWN5 [40] and PWLDA generalized gradient approximation (GGA): BP86, BLYP, PW91 [41, 42], PBE [43], mPWPW, mPWLYP hybrid GGA: B3LYP [44], PBE0 [45], X3LYP [46], mPW1PW, mPW1LYP, BHandHLYP [47] meta-GGA: TPSS [48], TPSSh [48], TPSS0 [48], M06L [49] hybrid meta-GGA: M062X [50] range-separated hybrids: $\omega$B97, [51] $\omega$B97X, [51]$\omega$B97X-D3, CAM-B3LYP [52], LC-BLYP [53], LC-PBE [54] double hybrid: $\omega$B97X-2 [55], B2PLYP [56], mPW2PLYP [57], B2GP-PLYP [58], PWPB95 [23], PBE-QIDH [59], PBE0-DH [60], $\omega$B2PLYP [61], $\omega$B2GP-PLYP [61], $\omega$B88PP86 [62], $\omega$PBEPP86 [62], SCS/SOS-B2PLYP21 [62], SCS-PBE-QIDH [62], SCS-B2GP-PLYP21 [62], SOS-B2GP-PLYP21 [62], SCS-$\omega$B2GP-PLYP21 [62], SOS-$\omega$B2GP-PLYP21 [62], SCS-RSX-QIDH [62], SOS-RSX-QIDH [62], SCS-$\omega$B88PP86 [62], SOS-$\omega$B88PP86 [62], SCS-$\omega$PBEPP86 [62], SOS-$\omega$PBEPP86 [62], DSD-BLYP [63], DSD-PBEP86 [63], DSD-PBEB95 [63], RSX-QIDH [64], and RSX-0H [65]. All DFT calculations were performed using Orca (version 6.0.0) [66, 67].
3 Bias Correction for Low-Variance Predictors
We examine the conditions under which regression-based bias correction is effective, to correct a low-variance, high-bias predictor using a low-bias, high-variance predictor. We clarify when and how the transformation preserves variance while ensuring bias reduction.
Definitions:
We begin by defining the key error metrics and terms used in this study.
  1. 1.
    Let $\hat{y}_{1}$ and $\hat{y}_{2}$ denote two predictive models (or predictors) for a given property. In this study, they represent two parameterizations of a DFA used for determining STG. Further, $\hat{y}_{\rm true}$ represents the exact theoretical value of the property under consideration. The best available finite-data estimate of $\hat{y}_{\rm true}$ is denoted as the reference value, $\hat{y}^{\rm ref}$. The $k$-th prediction from $\hat{y}_{1}$ and the corresponding reference value are denoted as $\hat{y}_{1,k}$ and $\hat{y}_{{\rm ref},k}$, respectively.
  2. 2.
    The expected value (i.e., mean) of $\hat{y}_{1}$ is given by $\mathbb{E}[\hat{y}_{1}]$. The variance of $\hat{y}_{1}$ around its mean is defined as: ${\rm Var}(\hat{y}_{1})=\mathbb{E}[\hat{y}_{1}^{2}]-\mathbb{E}[\hat{y}_{1}]^{2}$. The covariance between two predictors measures their linear relationship and is defined as: ${\rm Cov}(\hat{y}_{1},\hat{y}_{2})=\mathbb{E}[\hat{y}_{1}\hat{y}_{2}]-\mathbb{E}[\hat{y}_{1}]\mathbb{E}[\hat{y}_{2}]$.
  3. 3.
    The error metrics are defined as follows. Mean signed error (MSE): ${\rm MSE}=\mathbb{E}[\hat{y}_{1}-\hat{y}_{\rm ref}]$, mean absolute deviation (MAD): ${\rm MAD}=\mathbb{E}[|\hat{y}_{1}-\hat{y}_{\rm ref}|]$, standard deviation of error (SDE) for a sample of size $N$: ${\rm SDE}(\hat{y}_{1})=\sqrt{\frac{1}{N-1}\sum_{k=1}^{N}(\hat{y}_{1,k}-\hat{y}_{{\rm ref},k})^{2}}$. In the limit of a large sample (i.e., for a population), the square of SDE corresponds to the variance.
  4. 4.
    The term ‘bias’ refers to a systematic shift in a distribution and requires a reference for comparison. Throughout this study, bias specifically refers to systematic shifts in prediction error: $\hat{y}_{1}-\hat{y}_{\rm ref}$. Similarly, we distinguish between: ${\rm Var}(\hat{y}_{1})$ and ${\rm Var}(\hat{y}_{1}-\hat{y}_{\rm ref})$, where the former indicates spread around the mean while latter quantifies the spread of prediction errors, which is the primary focus of this study.
Bias Correction via Ordinary Least Squares Regression:
The bias correction scheme applies in cases where two predictors, $\hat{y}_{1}$ and $\hat{y}_{2}$, are available with known bias and variance relative to the reference values, $\hat{y}^{\rm ref}$. We assume that $\hat{y}_{1}$ has low variance in prediction errors (i.e., small SDE) but high bias (i.e., large magnitude of MSE), and $\hat{y}_{2}$ has low bias but high variance in prediction errors. These conditions are expressed as:
$|\mathbb{E}\left[\hat{y}_{1}-\hat{y}^{\rm ref}\right]|>0,\,\text{Var}\left(\hat{y}_{1}-\hat{y}^{\rm ref}\right)\approx 0;\qquad\mathbb{E}\left[\hat{y}_{2}-\hat{y}^{\rm ref}\right]\approx 0,\,\text{Var}\left(\hat{y}_{2}-\hat{y}^{\rm ref}\right)>0.$ (3.1)
We define a new estimator $\hat{y}_{1}^{*}$ by linearly scaling $\hat{y}_{1}$:
$\hat{y}_{1}^{*}$ $=$ $a\hat{y}_{1}+b,$ (3.2)
where $a$ and $b$ are determined via ordinary least squares regression, minimizing the squared error with respect to $\hat{y}_{2}$:
$\min_{a,b}\sum_{i}\left(a\hat{y}_{1,i}+b-\hat{y}_{2,i}\right)^{2}.$ (3.3)
The optimal values of $a$ and $b$ are given by:
$a=\frac{\text{Cov}(\hat{y}_{1},\hat{y}_{2})}{\text{Var}(\hat{y}_{1})};\quad b=\mathbb{E}[\hat{y}_{2}]-a\mathbb{E}[\hat{y}_{1}].$ (3.4)
Error analysis:
We want to determine whether the bias and variance of $\hat{y}_{1}^{*}$ match those of $\hat{y}_{2}$ and $\hat{y}_{1}$, respectively. The mean of $\hat{y}_{1}^{*}$ is:
$\mathbb{E}[\hat{y}_{1}^{*}]=\mathbb{E}[a\hat{y}_{1}+b]=a\mathbb{E}[\hat{y}_{1}]+b.$ (3.5)
Substituting the expression for $b$ from Eq. 3.4, we find:
$\mathbb{E}[\hat{y}_{1}^{*}]=a\mathbb{E}[\hat{y}_{1}]+\mathbb{E}[\hat{y}_{2}]-a\mathbb{E}[\hat{y}_{1}]=\mathbb{E}[\hat{y}_{2}].$ (3.6)
Thus, least-squares regression guarantees that the mean of $\hat{y}_{1}^{*}$ aligns with that of the reference—low-bias predictor $\hat{y}_{2}$—ensuring that the bias is corrected. The MSE of $\hat{y}_{1}^{*}$ follows:
$\mathbb{E}[\hat{y}_{1}^{*}-\hat{y}^{\rm ref}]$ $=$ $\mathbb{E}[a\hat{y}_{1}+b-\hat{y}^{\rm ref}]=\mathbb{E}[a\hat{y}_{1}]+\mathbb{E}[b]-\mathbb{E}[\hat{y}^{\rm ref}]$ (3.7)
$=$ $\mathbb{E}[a\hat{y}_{1}]+\left(\mathbb{E}[\hat{y}_{2}]-a\mathbb{E}[\hat{y}_{1}]\right)-\mathbb{E}[\hat{y}^{\rm ref}]$
$=$ $\mathbb{E}[\hat{y}_{2}]-\mathbb{E}[\hat{y}^{\rm ref}]=\mathbb{E}[\hat{y}_{2}-\hat{y}^{\rm ref}].$
Hence, the MSE of $\hat{y}_{1}^{*}$ is identical to that of $\hat{y}_{2}$, confirming that bias is fully transferred from $\hat{y}_{2}$.
The variance of $\hat{y}_{1}^{*}$ follows:
$\text{Var}(\hat{y}_{1}^{*})=\text{Var}(a\hat{y}_{1}+b)=a^{2}\text{Var}(\hat{y}_{1}).$ (3.8)
For $\hat{y}_{1}^{*}$ to preserve the variance of $\hat{y}_{1}$, we require $a^{2}=1$, which implies $a=\pm 1$. For $a\approx 1$, $\hat{y}_{1}^{*}$ retains the variance of $\hat{y}_{1}$ while aligning its mean with $\mathbb{E}[\hat{y}_{2}]$, effectively shifting the prediction distribution. However, in most practical cases, $|a|\neq 1$, leading to: $|a|<1$ resulting in $\text{Var}(\hat{y}_{1}^{*})<\text{Var}(\hat{y}_{1})$, or $|a|>1$ resulting in $\text{Var}(\hat{y}_{1}^{*})>\text{Var}(\hat{y}_{1})$. The variance of $\hat{y}_{1}^{*}-\hat{y}^{\rm ref}$ is given by:
${\rm Var}\left(\hat{y}_{1}^{*}-\hat{y}^{\rm ref}\right)$ $=$ ${\rm Var}\left(a\hat{y}_{1}+b-\hat{y}^{\rm ref}\right).$ (3.9)
Applying the variance sum formula, ${\rm Var}\left(A+B\right)={\rm Var}\left(A\right)+{\rm Var}\left(B\right)+{\rm Cov}\left(A,B\right)$, we arrive at:
${\rm Var}\left(\hat{y}_{1}^{*}-\hat{y}^{\rm ref}\right)$ $=$ ${\rm Var}\left(a\hat{y}_{1}\right)+{\rm Var}\left(b-\hat{y}^{\rm ref}\right)-2{\rm Cov}\left(a\hat{y}_{1},b-\hat{y}^{\rm ref}\right)$ (3.10)
$=$ ${\rm Var}\left(a\hat{y}_{1}\right)+{\rm Var}\left(\hat{y}^{\rm ref}\right)-2{\rm Cov}\left(a\hat{y}_{1},\hat{y}^{\rm ref}\right)$
$=$ ${\rm Var}\left(a\hat{y}_{1}-\hat{y}^{\rm ref}\right)$
$=$ $a^{2}{\rm Var}\left(\hat{y}_{1}-\hat{y}^{\rm ref}/a\right).$
This implies that the variance of $\hat{y}_{1}^{*}-\hat{y}^{\rm ref}$ is not simply $a^{2}$ times the variance of $\hat{y}_{1}-\hat{y}^{\rm ref}$, but is influenced by the distribution of the reference itself. Thus, only when $a=1$, the variance of $\hat{y}_{1}^{*}-\hat{y}^{\rm ref}$ exactly matches that of $\hat{y}_{1}-\hat{y}^{\rm ref}$. If $a=-1$, while the spread of $\hat{y}_{1}^{*}$ remains unchanged from the spread of $\hat{y}_{1}$, the spread of errors will increase. For general values of $a$, it is possible that fortuitously, the variance of prediction errors decreases after scaling. .
Practical Considerations:
In the above discussion, we designated the predictor with low MSE as an internal reference predictor, $\hat{y}_{2}$, interpreting it as the lower-bias model. However, this predictor will also have a low MAD only if all errors, $\hat{y}_{2}-\hat{y}^{\rm ref}$, have the same sign. In particular, if all errors are negative, then $\text{MAD}=-\text{MSE}$. In practice, a vanishing MSE for an approximate model often suggests a multimodal error distribution centered around zero, where positive and negative errors cancel each other out. In such cases, MSE can significantly underestimate the actual bias, making it an unreliable metric. A more robust measure of bias is MAD, which accounts for the magnitude of errors irrespective of sign. Thus, in this study, we select $\hat{y}_{2}$ as the predictor with the smallest MAD instead of the smallest MSE. This change does not affect any of the derivations presented above. Notably, the MSE of the scaled low-variance predictor, $\hat{y}_{1}^{*}$, remains close to that of $\hat{y}_{2}$, ensuring bias correction. Furthermore, since the variance of $\hat{y}_{1}^{*}$ remains approximately the same as that of $\hat{y}_{1}$ (for $a\approx 1$) the spread of errors is expected to be narrower. As a result, the MAD of $\hat{y}_{1}^{*}$ will typically be smaller than that of $\hat{y}_{2}$, reflecting a better balance between bias and variance.
4 Results and Discussions
Figure 2: Error metrics for various methods using the aug-cc-pVDZ basis set to predict 12 theoretical best estimates of STGs for triangular molecules. MAD, MSE, and SDE represent mean absolute deviation, mean signed error, and standard deviation, respectively. Methods are ordered by increasing MAD, with a line included as a visual guide. CC3 and EOM-CCSD results are from Ref [14].
4.1 Overall trends across methods
Figure 2 presents an overview of the performance of various theoretical methods using the TBE values from [14] as a reference. For consistency, all results are reported with the aug-cc-pVDZ basis set; results for other basis sets are provided in the Supplementary Information (SI), and the CC3 results are taken from [14]. Detailed error metrics for the wavefunction methods are provided in Table 1. For all methods listed in Table 1, increasing the basis set to aug-cc-pVTZ yields only minimal improvement, with EOM-CCSD being the sole exception, as it shows a slight deterioration in predictions with larger basis sets. The reference TBE S1 energies were computed at the CC3/aug-cc-pVTZ level with a post-CC3 correction via EOM-CCSDT (using the 6-31+G(d) basis set), while the T1 energies were obtained at the EOM-CCSD/aug-cc-pVTZ level with a post-CCSD correction via CC3 using the aug-cc-pVDZ basis set [14]. Although the TBE approach treats S1 and T1 energies using different protocols, both are estimated at higher levels of theory currently available.
Among all methods, CC3/aug-cc-pVDZ shows the lowest values for MSE, MAD, and SDE ($-1$, 10, and 13 meV (milli-eV), respectively; see Table 1). When combined with the aug-cc-pVTZ basis set, CC3 essentially reproduces the TBE results. The next best performance is observed for L-CC2 with an MAD and SDE of 16 and 13 meV. Overall, when using our more accurate value for system-10, ADC(2)/aug-cc-pVTZ yields an MAD and SDE of 20 and 16 meV, compared to 30 and 35 meV from Ref. [14]. Our L-ADC(2)/aug-cc-pVTZ results show an MAD and SDE of 22 and 13 meV, respectively, which are similar to the corresponding ADC(2) values of 20 and 16 meV.
Both the mean error and the spread of the error are significant for EOM-CCSD (Figure 2). Its MSE indicates that its predictions can serve as an upper bound for negative STGs; however, an SDE of approximately 80–90 meV diminishes confidence in predictions of gaps smaller than 0.1 eV, potentially leading to false negatives (i.e., systems with truly negative gaps might be predicted as having positive gaps). The large SDE suggests that the error is not systematic, so not all small, positive STGs predicted by EOM-CCSD can be reliably classified as negative. In particular, the similarity-transformed local variant, DLPNO-STEOM-CCSD, exhibits a bias exceeding 0.25 eV with an SDE around 0.1 eV (see Figure 2), rendering it unsuitable for applications involving negative STG systems [68, 69].
Table 1: Error metrics for predicting twelve values of S1 & T1 energies, and STG of triangular molecules. For various methods, values are reported compared to the theoretical best estimates from Ref. [14]. MSE: mean signed error, MAD: mean absolute deviation, SDE: standard deviation of the error, minE: minimal error, and maxE: maximal error. In all cases, the basis set is aug-cc-pVDZ. Wherever necessary, the convention for deviation is ‘Method--TBE.’ All values are in eV.
Method Energy MSE MAD SDE minE maxE
L-CC2/aug-cc-pVDZa S1 0.047 0.047 0.016 0.020 0.071
T1 0.060 0.060 0.017 0.038 0.091
STG --0.013 0.016 0.013 --0.028 0.010
L-CC2/aug-cc-pVTZa S1 0.051 0.051 0.010 0.035 0.064
T1 0.065 0.065 0.016 0.032 0.090
STG --0.014 0.015 0.010 --0.028 0.005
L-ADC(2)/aug-cc-pVDZa S1 0.006 0.031 0.035 --0.056 0.056
T1 0.025 0.030 0.032 --0.020 0.078
STG --0.019 0.023 0.017 --0.038 0.023
L-ADC(2)/aug-cc-pVTZa S1 0.003 0.017 0.021 --0.037 0.039
T1 0.023 0.023 0.018 --0.002 0.052
STG --0.020 0.022 0.013 --0.036 0.009
ADC2/aug-cc-pVTZa S1 -0.000 0.021 0.026 -0.051 0.050
T1 0.015 0.022 0.022 -0.023 0.049
STG -0.016 0.020 0.016 -0.039 0.021
CC3/aug-cc-pVDZb S1 0.001 0.015 0.018 --0.030 0.026
T1 0.002 0.016 0.020 --0.038 0.037
STG --0.001 0.010 0.013 --0.035 0.015
  • a This work
  • b From Ref. [14]
Among the DFT methods, the dh-DFT approaches B2GP-PLYP and PBE-QIDH display very small SDEs, indicating that their deviations from the TBE values are primarily systematic. The spin-component-scaled (SCS) and opposite-spin-scaled (SOS) variants of these methods exhibit reduced MSE and MAD at the expense of larger SDEs. With few exceptions, the remaining DFAs show large errors, rendering them unreliable for first-principles identification of negative STG systems.
Figure 3: Bias-variance tradeoff in various methods as a scatterplot of MAD versus SDE for predicting 12 theoretical best estimates of STGs in triangular molecules. Wavefunction methods are marked with blue stars, and DFAs with red circles. For clarity, names are shown only for selected points. All results are based on the aug-cc-pVDZ basis set. CC3 and EOM-CCSD results are from Ref [14]
4.2 Bias-variance tradeoff across methods
Since the typical magnitudes of negative STGs reported in various studies are $\approx 0.1$ eV or less, methods with mean errors exceeding 0.05 eV (50 meV) are generally unacceptable for drawing qualitative conclusions. Figure 2 reveals interesting trends: while some methods exhibit a high MAD, they maintain a very low SDE. These cases are highlighted in Figure 3, which plots MAD versus SDE for various methods (excluding those with SDE $>40$ meV or MAD $>250$ meV). Overall, CC3 and L-CC2 demonstrate superior performance with low MAD and SDE, whereas several DFT approaches—notably, B2GP-PLYP and PBE-QIDH—exhibit remarkably small variance. In particular, PBE-QIDH shows an SDE of less than 10 meV, indicating that its predictions are narrowly clustered around the TBE values but shifted by a constant systematic error. Although the MAD of PBE-QIDH exceeds 175 meV, systematic errors are generally easier to correct through linear regression than the non-systematic errors associated with high variance. For properties such as atomization or formation energies, DFT predictions are often corrected for systematic errors via dressed atom corrections, leading to lower errors for DFAs on the higher rungs of Jacob’s ladder [70, 56, 71].
Among the local correlated methods, L-CC2 exhibits MAD and SDE values close to those of CC3, while L-ADC2 shows a slightly larger SDE than L-CC2 (Figure 3). In contrast, the dh-DFT methods B2GP-PLYP and PBE-QIDH display SDEs that are smaller than L-CC2’s, but their MADs are larger by about an order of magnitude. To explore the potential of low-variance, high-bias methods, we applied a least-squares correction to the STG predictions of B2GP-PLYP, PBE-QIDH, L-ADC(2), and L-CC2 using TBE values as the reference; the results are collected in Table 2. As expected, the MSE becomes zero after linear correction. L-ADC(2) and L-CC2 show MADs of 0.023 and 0.016 eV with SDEs of 0.017 and 0.013 eV, respectively, with only slight improvement upon correction, indicating that their errors are primarily non-systematic. In contrast, dh-DFT methods B2GP-PLYP and PBE-QIDH, which initially exhibit MADs of 0.119 and 0.177 eV and SDEs of 0.011 and 0.009 eV respectively, experience a dramatic drop in MAD (to below 0.01 eV) after correction, while retaining their low SDEs. This implies that the residual errors in these dh-DFT methods are predominantly systematic. A similar analysis for the SCS and SOS variants of these methods, presented in Table 3, confirms that all fixed dh-DFT methods show reduced MAD and SDE values, indicating that their prediction errors are largely systematic when compared to TBE.
Table 2: For various methods (aug-cc-pVDZ basis set) STGs of benchmark systems (Figure 1) are presented before and after bias-correction. Error metrics and the reference TBE values are also provided. All values are in eV.
# B2GP-PLYP PBE-QIDH L-ADC2 L-CC2 TBE
corr.a corr.b corr.c corr.d
1 --0.095 --0.220 --0.035 --0.215 --0.255 --0.229 --0.247 --0.229 --0.219
2 --0.007 --0.126 0.037 --0.141 --0.134 --0.116 --0.128 --0.116 --0.131
3 0.020 --0.098 0.081 --0.095 --0.120 --0.103 --0.109 --0.098 --0.101
4 --0.014 --0.134 0.062 --0.115 --0.132 --0.114 --0.130 --0.118 --0.119
5 0.005 --0.114 0.082 --0.094 --0.138 --0.119 --0.127 --0.115 --0.103
6 0.042 --0.074 0.094 --0.082 --0.102 --0.086 --0.086 --0.076 --0.071
7 0.070 --0.044 0.122 --0.053 --0.080 --0.065 --0.069 --0.059 --0.042
8 0.081 --0.033 0.132 --0.043 --0.063 --0.049 --0.052 --0.043 --0.029
9 --0.080 --0.204 --0.017 --0.196 --0.220 --0.197 --0.215 --0.199 --0.199
10 --0.173 --0.303 --0.126 --0.308 --0.318 --0.289 --0.325 --0.304 --0.305
11 --0.049 --0.171 --0.005 --0.184 --0.199 --0.177 --0.187 --0.172 --0.195
12 0.177 0.070 0.246 0.074 0.085 0.090 0.072 0.075 0.062
MSE 0.119 0.000 0.177 $0$-0.019 $0$-0.013 0.000
MAD 0.119 0.007 0.177 0.008 0.023 0.014 0.016 0.010
SDE 0.011 0.009 0.009 0.009 0.017 0.016 0.013 0.012
  • a Corrected with ${\rm slope}=1.0654$ and ${\rm intercept}=-0.1190$
  • b Corrected with ${\rm slope}=1.0284$ and ${\rm intercept}=-0.1787$
  • c Corrected with ${\rm slope}=0.9409$ and ${\rm intercept}=-0.0103$
  • d Corrected with ${\rm slope}=0.9546$ and ${\rm intercept}=-0.0065$
  • e From Ref. [14]
Table 3: Error metrics for predicting twelve values of S1 & T1 energetics of triangular molecules reported in Ref. [14]. Values are reported compared to the theoretical best estimates (TBE) from Ref. [14]. In all cases, the basis set is aug-cc-pVDZ, and the geometries are from Ref. [14]. MSE: mean signed error, MAE: mean absolute error, SDE: standard deviation of the error, minE: minimal error, and maxE: maximal error. Wherever necessary, the convention for deviation is ‘Method--TBE.’ All values are in eV.
Method Energy Slope Intercept Before correction After correction
MAE SDE MAE SDE
B2GP-PLYP S1 1.0350 -0.2382 0.184 0.030 0.012 0.020
T1 1.0294 -0.1105 0.065 0.031 0.018 0.024
STG 1.0654 -0.1190 0.119 0.011 0.007 0.009
SOS-B2GP-PLYP21 S1 1.0164 -0.2574 0.231 0.015 0.008 0.011
T1 1.0048 -0.1889 0.181 0.019 0.013 0.018
STG 0.9167 -0.0561 0.050 0.013 0.007 0.010
SCS-B2GP-PLYP21 S1 1.0086 -0.2948 0.281 0.013 0.009 0.012
T1 1.$0$0.2164 0.216 0.020 0.015 0.020
STG 0.9246 -0.0685 0.064 0.012 0.008 0.009
PBE-QIDH S1 0.9627 -0.2158 0.277 0.034 0.013 0.023
T1 0.9693 -0.0514 0.100 0.033 0.017 0.025
STG 1.0284 -0.1787 0.177 0.009 0.008 0.009
SOS-PBE-QIDH S1 1.0120 -0.1281 0.110 0.012 0.008 0.009
T1 0.9889 -0.0862 0.104 0.017 0.012 0.016
STG 0.8423 -0.0246 0.018 0.021 0.010 0.012
SCS-PBE-QIDH S1 1.0091 -0.1479 0.134 0.011 0.007 0.009
T1 0.9907 -0.0818 0.096 0.018 0.013 0.017
STG 0.8815 -0.0477 0.038 0.017 0.009 0.011
  • a Calculated with STG calculated using linearly corrected S1 and T1 values.
  • b Calculated using linearly corrected STG values.
4.3 Internally referenced bias-variance correction for double-hybrid DFT
While one can use L-CC2, CC3, or TBE values to correct PBE-QIDH predictions through linear scaling, we further investigated whether the low-variance DFT methods—B2GP-PLYP and PBE-QIDH—can be adjusted to minimize bias. The PBE-QIDH DFA is defined as follows [59]:
$E_{\rm xc}^{\rm PBE-QIDH}\left[\rho\right]$ $=$ $a_{x}E_{x}^{\rm HF}+\left(1-a_{x}\right)E_{x}^{\rm PBE}+a_{c}E_{c}^{\rm MP2}+\left(1-a_{c}\right)E_{c}^{\rm PBE}.$ (4.1)
Here, the coefficients $a_{x}$ and $a_{c}$ control the fraction of exact exchange (via Hartree–Fock) and MP2-level correlation, respectively. In the standard formulation, PBE-QIDH employs $a_{x}=0.69$ and $a_{c}=0.33$ [72] and belongs to a family of dh-DFAs with $a_{c}=a_{x}^{3}$, which have been found to deliver low errors for STG [73]. Similarly, B2GP-PLYP is based on the GGA-B88 exchange functional and a combination of LYP-GGA and VWN3-LDA correlation functionals:
$E_{\rm xc}^{\rm B2GP-PLYP}\left[\rho\right]$ $=$ $a_{x}E_{x}^{\rm HF}+\left(1-a_{x}\right)E_{x}^{\rm B88}+a_{c}E_{c}^{\rm MP2}+\left(1-a_{c}\right)\left(E_{c}^{\rm VWN3}+E_{c}^{\rm LYP}\right)$ (4.2)
with standard mixing coefficients $a_{x}=0.65$ and $a_{c}=0.36$ [74, 75, 76].
Figure 4: Dependence of prediction errors of dh-DFT with exchange and correlation mixing parameters $a_{x}$ and $a_{c}$ (See Eq. 4.1 and Eq. 4.2). MSE, MAD and SDE (in meV) are shown B2GP-PLYP (left panels) and PBE-QIDH (right panels) compared to TBE. In all plots, the default mixing coefficients are marked by a red cross—(0.65, 0.36) for B2GP-PLYP and (0.69, 0.33) for PBE-QIDH—and coefficients that minimize MAD are marked by a blue star—(0.75, 0.55) for both B2GP-PLYP and PBE-QIDH.
Prior studies have shown that the $a_{x}$-$a_{c}$ parameter space of B2GP-PLYP exhibits a continuous range of optimal parameters depending on the target property [58, 23]. For instance, Tarnopolsky et al. demonstrated variations in the SDE for thermochemistry and reaction barriers as a function of $a_{x}$ and $a_{c}$, and proposed the B2T-PLYP and B2K-PLYP df-DFAs with ($a_{x}=0.60$, $a_{c}=0.31$) and ($a_{x}=0.72$, $a_{c}=0.40$), respectively [22]. However, while these studies focused on a single error metric across different properties—emphasizing variance because bias in thermochemistry is corrected through quasi-atomic corrections[70, 56, 71]—the bias-variance trade-off for a single property such as the STG has not been thoroughly examined.
Using B2GP-PLYP and PBE-QIDH DFAs, we scanned the $a_{x}$ and $a_{c}$ values and computed error metrics for all 12 benchmark STGs using TBE as the reference. Some calculations failed to converge when $a_{x}$ and $a_{c}\approx 0$. As shown in Figure 4, both dh-DFAs exhibit low SDEs for the default mixing parameters, and a continuous range of ($a_{x}$, $a_{c}$) combinations yielding zero MSE is observed (Figure 4, top panels). However, the minimum MAD is achieved at $a_{x}=0.75$ and $a_{c}=0.55$, with both DFAs delivering MADs of approximately 25 meV compared to TBE. We denote these optimized parameterizations as B2GP-PLYP (75,55) and PBE-QIDH (75,55). To assess the variation of optimal $a_{x}$ and $a_{c}$ across error metrics for S1 and T1 energies separately, we performed a comparative analysis (Table S1) using PBE-QIDH. Notably, increased accuracy in S1 energies is observed when both mixing parameters approach 1 (with $a_{x}=0.90$ and $a_{c}=0.85$ yielding a MAD of 68 meV), while the lowest MAD for T1 (22 meV) is obtained for $a_{x}=0.5$ and $a_{c}=0.15$.
To apply the scaling protocol discussed in Section 3, the default B2GP-PLYP and PBE-QIDH parameterizations serve as the low-variance estimators ($y_{1}$ in Eq.(3.1)), while the (75,55) variants, which exhibit the lowest MAD, serve as the reference low-bias estimators ($y_{2}$ in Eq.(3.1)). We then determine the slope and intercept via linear regression to correct the bias, with the corrected predictions denoted as B2GP-PLYP and PBE-QIDH. Table 4 presents the STG values for the 12 benchmark systems predicted by the bias-corrected dh-DFAs. Importantly, the scaling procedure does not require any additional high-level reference data to identify the low-bias or low-variance domain in the parameter space of dh-DFAs. The bias-corrected predictions, which simultaneously reflect the small SDE of the base dh-DFAs and the low MAD of the internal references, achieve MAD||||SDE values of 10||||12 and 9||||11 meV for B2GP-PLYP and PBE-QIDH, respectively, with respect to TBE.
Table 4: STGs of benchmark systems shown in Figure 1 calculated with double-hybrid DFT methods using the aug-cc-pVDZ basis set. The error metrics MSE (mean signed error), MAD (mean absolute deviation), and SDE (standard deviation of the error) are with respect to theoretical best estimates given in Table 2. All values are in eV.
# B2GP-PLYP PBE-QIDH
(65, 36) (75, 55) scaleda (69, 33) (75, 55) scaledb
1 --0.095 --0.245 --0.231 --0.035 --0.239 --0.216
2 --0.007 --0.102 --0.130 0.037 --0.094 --0.138
3 0.020 --0.104 --0.099 0.081 --0.096 --0.089
4 --0.014 --0.178 --0.138 0.062 --0.170 --0.110
5 0.005 --0.153 --0.116 0.082 --0.144 --0.088
6 0.042 --0.064 --0.074 0.094 --0.056 --0.075
7 0.070 --0.034 --0.041 0.122 --0.024 --0.045
8 0.081 --0.016 --0.029 0.132 --0.007 --0.034
9 --0.080 --0.227 --0.214 --0.017 --0.222 --0.197
10 --0.173 --0.299 --0.321 --0.126 --0.295 --0.316
11 --0.049 --0.154 --0.178 --0.005 --0.148 --0.183
12 0.177 0.086 0.082 0.246 0.095 0.091
MSE 0.119 --0.003 --0.003 0.177 0.004 0.004
MAD 0.119 0.025 0.010 0.177 0.027 0.009
SDE 0.011 0.030 0.012 0.009 0.030 0.011
  • a Calculated by scaling B2GP-PLYP (65,36) values using slope = 1.1521 and intercept=$-0.1220$
  • b Calculated by scaling PBE-QIDH (69,33) values using slope = 1.0940 and intercept=$-0.1780$
Figure 5: Comparison of 12 STGs of benchmark systems across methods: a) Scatterplot of values predicted by various methods (with the aug-cc-pVDZ basis set) versus theoretical best estimates (TBE). b) Distribution of errors in predicted STGs relative to TBE shown as smooth densities.
The overall correlation between the predicted values and TBE is illustrated in Figure 5. Due to the large bias in the original dh-DFT predictions, the Pearson correlation coefficients are strongly negative ($\rho=-0.61$ for B2GP-PLYP and $\rho=-2.45$ for PBE-QIDH; see Figure 5a). Upon applying the internally referenced scaling, the correlation improves markedly to $\rho=0.98$ for both DFAs, with all 12 values aligning with TBE more closely than those from L-ADC(2), L-CC2, or even CC3. The overall performance of the scaling procedure stems from the exceptionally small variance of the dh-DFAs and the low MAD achieved with the (75,55) parameterizations.
4.4 Application to heptazine derivatives
To probe the applicability of our bias-correction strategy, we focused on four DFIST candidates proposed in Ref. [4]. At the T1 geometry, the STGs of HzTFEX2, HzPipX2, HzTFEP2, and HzTFET2 at the EOM-CCSD, ADC(2), L-CC2, and SCS-ADC(2) levels exhibit smaller magnitudes[4] than those of the twelve triangular systems discussed above. This difference is attributed to the change in geometry, as STGs generally increase when moving away from the ground state minimum (S0) as pointed out from the context of excited-state nuclear dynamics[77] and the pseudo-Jahn–Teller effect [25]. Despite the larger molecular size of the heptazine derivatives compared to the triangular systems, local methods such as L-ADC(2) and L-CC2 provide significant speedups over their canonical counterparts, while dh-DFT calculations remain computationally feasible. This enables a straightforward application of our internally referenced scaling procedure, which corrects systematic errors without requiring external high-level reference data.
Figure 6: Heptazine derivatives proposed as candidates for light emitters exhibiting delayed fluorescence from inverted singlet and triplet excited states (DFIST) in Ref. [4] are shown along with their computed STG values (in eV) from various methods. B2GP-PLYP and PBE-QIDH indicate that the results of the base methods B2GP-PLYP and PBE-QIDH are corrected with respect to their variants with $a_{x}=0.75$ and $a_{c}=0.55$; more details are available in the SI. Atoms in white||||black||||blue||||red||||cyan are H||||C||||N||||O||||F.
For all four systems, geometry optimization was performed at the $\omega$B97X-D3/def2TZVP level (see SI for details), followed by single-point excited-state energy calculations. The STGs of these four systems, obtained from various methods, are presented in Figure 6. In the absence of a high-fidelity reference such as theoretical best estimates for these systems, L-CC2 results are considered the most reliable. We used B2GP-PLYP (75,55) energies as references to scale the predictions from B2GP-PLYP (65,36) and similarly corrected PBE-QIDH (69,33) predictions using the (75,55) configuration. Notably, the application of the internal scaling does not require any additional high-level reference to identify the low-bias or low-variance domain in the parameter space of dh-DFAs.
Among the results from various dh-DFAs, SCS-PBE-QIDH shows the best agreement with L-CC2, achieving a MAD of 0.014 eV. In contrast, bias-corrected results for B2GP-PLYP and PBE-QIDH are systematically more negative than the L-CC2 values, with MADs of 0.057 eV and 0.048 eV, respectively. While one might argue that simply shifting the results of these two DFAs could render the STGs negative, our correction is performed without requiring reference values from another method. As in the case of triangulene systems, B2GP-PLYP exhibits a smaller bias compared to PBE-QIDH, as reflected in its predicted STGs of --0.132, --0.042, --0.107, and --0.105 eV. Furthermore, we observed that L-CC2’s error relative to the TBE is consistently shifted by -0.013 eV, whereas the errors of the (75,55) variant of the two dh-DFAs are centered around zero (albeit with a large spread), as shown in Figure 5. Consequently, compared to L-CC2, the bias-corrected dh-DFAs are expected to exhibit a non-vanishing MSE; indeed, the MSE values for B2GP-PLYP (75,55) and PBE-QIDH (75,55) are -0.057 eV and -0.048 eV, respectively. Compared to a more accurate reference such as CC3 (which is practically challenging for the four heptazine derivatives), our bias-corrected results may show better agreement than the other methods discussed. Moreover, it is possible to identify combinations of $a_{x}$ and $a_{c}$ that yield low bias and low variance relative to L-CC2, which can be further explored for extending the bias correction scheme to larger datasets lacking high-level reference data.
5 Conclusions
First principles excited state modeling plays a crucial rule in the identification of new molecular light emitters with a negative S1-T1 energy gap. Historically, theoretical interest in such molecules existed for a long time in the context of Hund’s rule violation[78, 79, 80, 81, 82]. Yet, revived interest in them increased since their identification through first principles modeling of heptazine[83, 84]. Although individual S1 and T1 energies are challenging to predict with high precision, the error cancellation inherent in STG calculations allows for improved accuracy, even when the reported gaps are less than 0.1 eV. Benchmarking twelve STG energy gaps in triangular molecules, we find that local wavefunction methods—specifically L-CC2 and L-ADC2—deliver mean errors of approximately 15 meV with correspondingly low standard deviations, providing both exceptional accuracy and the computational speed required for high-throughput screening of medium size molecules such as azaphenalenes with bulky substitutions.
In parallel, we show that double-hybrid DFT approximations—though initially marked by high systematic errors—can be significantly improved through linear regression-based bias correction without depending on a high-level reference. For example, the PBE-QIDH method in its default configuration (69% exact exchange and 33% MP2-level correlation) exhibits an impressively low variance (9 meV) but high mean errors; by adjusting the parameters to 75% exchange and 55% correlation, the mean error is reduced to 4 meV, albeit with increased variance. When applying this strategy to four heptazine derivatives, the SCS and SOS variants of both B2GP-PLYP and PBE-QIDH—though not subjected to bias correction—yield STG predictions in better agreement with L-CC2, suggesting that these variants naturally mitigate some of the bias in standard dh-DFT. However, bias correction remains valuable for further refining low-variance models, particularly when an appropriate low-bias reference is available.
The purpose of this study is not to advocate different models for different properties but to leverage the intrinsic bias-variance tradeoff for accurate STG predictions. By using a suitable low-bias method as an internal reference, our approach corrects systematic errors in low-variance models, thereby enabling quantitatively reliable negative STG estimations. Open questions remain regarding the performance of these parameterized models for positive STGs and systems with charge-transfer excitations[85], where even methods like CC2 may exhibit deficiencies. Robust application of our strategy requires reliable reference datasets and could benefit from further bias-variance analysis using alternative information-based metrics [86]. Overall, our findings establish a mathematically sound framework for predictive modeling in first-principles molecular design, with promising implications for the development of novel light emitters and other applications.
6 Supplementary Information
i) Dataset of S1, T1 and S1-T1 energies of twelve triangular molecules calculated with various methods and basis sets. ii) Figures S1–S3 present error metrics for cc-pVDZ, cc-pVTZ, and aug-cc-pVTZ basis sets; iii) Tables S1 presents optimal $a_{x}$ and $a_{c}$ and the associated error metrics for different energies; iv) Tables S2–S3 present excited state energies of heptazine derivatives; v) Equilibrium coordinates of heptazine derivatives.
7 Data Availability
The data that support the findings of this study are within the article and its supplementary material.
8 Acknowledgments
We thank Prof. Denis Jacquemin for commenting on the ADC(2) and CC2 results of the benchmark triangular molecules. We acknowledge the support of the Department of Atomic Energy, Government of India, under Project Identification No. RTI 4007. All calculations have been performed using the Helios computer cluster, which is an integral part of the MolDis Big Data facility, TIFR Hyderabad (http://moldis.tifrh.res.in).
9 Author Declarations
9.1 Author contributions
AM: Conceptualization (equal); Analysis (equal); Data collection (equal); Writing (equal). RR: Conceptualization (equal); Analysis (equal); Data collection (equal); Funding acquisition; Project administration and supervision; Resources; Writing (equal).
9.2 Conflicts of Interest
The authors have no conflicts of interest to disclose.